# Vision Transformers for Multi-Variable Climate Downscaling: Emulating Regional Climate Models with a Shared Encoder and Multi-Decoder Architecture

### v1

@@ Abstract
Global Climate Models (GCMs) are critical for simulating large-scale climate dynamics, but their coarse spatial resolution limits their applicability in regional studies. Regional Climate Models (RCMs) refine this through dynamic downscaling, albeit at considerable computational cost and with limited flexibility. While deep learning has emerged as an efficient data-driven alternative, most existing studies have focused on single-variable models that downscale one variable at a time. This approach can lead to limited contextual awareness, redundant computation, and lack of cross-variable interaction. Our study addresses these limitations by proposing a multi-task, multi-variable Vision Transformer (ViT) architecture with a shared encoder and variable-specific decoders (1EMD). The proposed architecture jointly predicts three key climate variables: surface temperature (tas), wind speed (sfcWind), and 500 hPa geopotential height (zg500), directly from GCM-resolution inputs, emulating RCM-scale downscaling over Europe. We show that our multi-variable approach achieves positive cross-variable knowledge transfer and consistently outperforms single-variable baselines trained under identical conditions, while also improving computational efficiency. These results demonstrate the effectiveness of multi-variable modeling for high-resolution climate downscaling.
@@ 1 Introduction
Global Climate Models (GCMs) are essential tools for simulating the Earth’s climate system and projecting future climate change. However, their coarse spatial resolution, typically around 1°, limits their ability to accurately capture regional-scale dynamics. To overcome this limitation, Regional Climate Models (RCMs) are employed to dynamically downscale GCM outputs, providing finer-scale climate information at resolutions typically ranging from 10 to 50 kilometers. Although RCMs enhance spatial detail and realism, they are computationally demanding, limiting the number of possible model runs in the context of regional coordinated experiments [1].
@@
An emerging alternative to traditional numerical downscaling involves the use of deep learning techniques to emulate the RCM downscaling process. Instead of performing computationally intensive simulations, deep learning models learn to reproduce the mapping from coarse-resolution GCM outputs to high-resolution RCM fields from available data. This approach presents a rapid, data-driven method for producing detailed climate projections applicable across multiple regions and scenarios, increasing the number of available simulations [2].
@@
In this study, we frame the GCM-to-RCM emulation task as a supervised deep learning problem. Using paired datasets from GCM and RCM outputs, we employ Vision Transformer (ViT) models to learn the relationship between coarse-resolution inputs and their high-resolution counterparts over the European domain. The primary objective of this approch is to produce spatially detailed outputs that accurately capture the fine-scale structures inherent to RCM simulations, enabling efficient and cost-effective generation of high-resolution climate data.
@@
The main contribution of this work is the exploration of joint, multi-variable processing. Typically, downscaling meteorological variables from GCM to RCM involves handling numerous diverse variables spanning atmospheric, surface, and radiation categories. Most existing deep learning approaches for climate downscaling focus exclusively on single-variable modeling, training and applying uni-variate models separately on each variable [3, 4, 5, 6, 7]. This approach poses several issues: single-variable models cannot exploit valuable cross-variable interactions, managing and executing multiple independent models is inefficient, and similar meteorological variables result in redundant learning efforts when modeled separately. In response, this study explores multi-variable processing strategies.
@@
Figure 1: Visualization of a single daily sample for the three studied variables: surface temperature (left column), wind speed (center column), and geopotential height (right column). Each variable is shown for both the high-resolution Regional Climate Model ICTP-RegCM4-6 (top row) and the low-resolution Global Climate Model MPI-ESM-LR (bottom row) re-gridded onto the same European domain.
@@
Vision Transformers (ViTs) have emerged as powerful architectures in image-to-image translation and super-resolution tasks due to their attention-based processing and ability to capture long-range dependencies [8, 9, 2, 10]. Leveraging this architecture, we propose processing several meteorological variables simultaneously, enabling the effective exploitation of inter-variable relationships and potentially enhancing overall performance and physical coherence. This approach significantly reduces the number of required models, thereby decreasing computational demands per variable and improving efficiency. However, multi-variable modeling also introduces challenges such as potential training instabilities and cross-variable interference, which may manifest as artifacts in outputs.
@@
To evaluate multi-variable modeling strategies for GCM-to-RCM downscaling, we leverage modern ViT architectures to model this image-to-image translation problem. Our study focuses on three common meteorological variables: wind speed (sfcWind), temperature (tas), and geopotential height at 500 hPa (zg500). An image depicting a sample of the three variables for both RCM and GCM is reported in Figure 1. We first establish baseline performance using single-variable ViT models trained independently for each variable. Subsequently, we introduce two multi-variable approaches: one treating variables as separate channels (1E1D), and a second employing a single shared encoder with multiple dedicated decoders (1EMD), one per variable. We ensure a robust and reproducible comparison by standardizing model structure, parameter count, training epochs, learning rates, and optimizers across all experiments.
@@
Our findings reveals that multi-variable modeling offers clear advantages in terms of computational efficiency. However, the simpler multi-channel approach (1E1D) suffers from cross-variable interference, failing to match the performance of single-variable baselines. In contrast, our structured multi-decoder design (1EMD) avoids these issues and outperforms single-variable models across all tested variables. These results highlight the effectiveness of our multi-variable architecture for high-resolution climate downscaling.
@@ 1.1 Data Selection
Our study is based on the EURO-CORDEX data. EURO-CORDEX, developed as part of the Coordinated Regional Downscaling Experiment (CORDEX) under the World Climate Research Programme (WCRP), provides a consistent set of high-resolution (0.11°) RCM simulations, driven by boundary conditions from multiple GCMs and emission scenarios. These simulations are widely used for regional climate impact assessments across Europe [1]. We subjectively selected one GCM-RCM couple from the available simulations: the MPI-ESM-LR GCM with the ICTP-RegCM4-6 RCM simulation under emissions scenario RCP8.5.
The selection of surface air temperature (tas), surface wind speed (sfcWind), and geopotential height at 500 hPa (zg500) was guided by a combination of practical relevance and methodological interest. The three variables exhibit markedly different spatial characteristics when downscaled. Surface temperature tends to produce relatively smooth spatial fields varying mostly with elevation and latitude, surface wind speed fields often feature localized and high-frequency patterns, and geopotential height exhibits broad, smooth spatial gradients that reflect the large-scale atmospheric circulation patterns in the mid-troposphere. This diversity offers a useful testbed for evaluating how well data-driven models handle different signal types, and whether modeling them jointly can facilitate positive transfer across variables.
The choice was also shaped by the broader goal of exploring multi-variable downscaling. By selecting a small but diverse set of variables, we can begin to assess whether downscaling performance benefits from shared representations across related but physically distinct fields. The downscaling task addresses a spatial resolution increase from 1.0° to 0.1°, representing a tenfold enhancement. This is among the largest resolution gaps available in publicly accessible GCM-RCM datasets, offering a particularly demanding test case for evaluating the capabilities of data-driven downscaling methods.
@@ 1.2 Related Works
In recent years, the focus of climate downscaling has increasingly shifted toward deep learning models [11, 12, 3]. A wide range of architectures has been explored to downscale variables such as temperature, wind components, surface pressure, and precipitation, among others. These include convolutional neural networks (CNNs) [12, 13, 14, 15, 16, 4], U-Nets [6, 17, 18], Vision Transformers (ViTs) [8, 9, 10], generative adversarial networks (GANs) [19, 20, 21, 22], and diffusion models [23, 24, 25, 26, 27].
Different architectures show varying strengths depending on the target variable. Generative models like GANs and diffusion models are particularly effective at preserving high-frequency spatial detail and have shown strong performance in downscaling wind speed [28]. In contrast, smoother fields such as temperature are often better handled by U-Nets [29]. Despite these findings, there is no clear consensus on a universally optimal architecture. Recently, ViT-based models have shown competitive and often superior performance across multiple climate variables, outperforming CNN-based baselines in several studies [30, 31, 8]. As such, ViTs form the foundation of our proposed approach.
ViT-based architectures for image-to-image tasks often draw inspiration from the encoder-decoder structure of U-Nets. One common design integrates a transformer-based encoder with a convolutional decoder [32, 33, 34], while other variants opt for fully transformer-based pipelines [35, 36]. A key architectural consideration is the treatment of skip connections, which preserve spatial resolution and improve gradient flow. These may be preserved directly [32, 37], omitted entirely under the assumption of rich encoder representations [33], or implemented through attention-based gating mechanisms [35].
@@
Multi-variable downscaling naturally aligns with the framework of multi-task learning, where a single model learns to solve multiple related tasks simultaneously. In climate applications, multi-task learning offers several potential advantages: improved generalization through shared representations, reduced training and inference cost, and the possibility of positive transfer across variables [38].
Several recent studies support the promise of this approach. For instance, Ling et al. [39] present a multi-task model for seasonal climate prediction that jointly forecasts several indices related to the Indian Ocean Dipole. The model is trained using a loss function defined as a weighted sum of the individual task with a higher loss weight to the primary target (DMI). The model achieves improved performance compared to single-variable baselines. Similarly, Fan et al. [40] propose M2GSNet, a graph-based multi-task model for wind power prediction across multiple farms. By utilizing shared encoders across tasks and training with an unweighted sum of L1 losses across farms, the model achieves superior forecasting accuracy compared to single-variable baselines.
Other examples further support the benefits of multi-task formulations. Bannai et al. [41] use a shared CNN backbone with two heads for simultaneous rain/no-rain classification and rain rate estimation, yielding improved performance through joint optimization of a combined loss function. Liu et al. [42] propose a two-stage downscaling system incorporating bias correction, terrain inputs, and a multi-task loss with physical constraints, demonstrating the potential of multi-variable models for enhancing both accuracy and realism in climate data products.
@@ 2 Methodology
Figure 2: Comparison between a single-variable encoder-decoder model and multi-variable models using either a single shared decoder (1E1D) or multiple decoders branching from a shared encoder (1EMD)
@@ 2.1 Architectural Backbone: Vision Transformer (ViT)
In recent years ViTs have been successfully applied to a variety of super-resolution tasks [43, 44, 45, 46]. Our proposed ViT architecture belongs to the ViT-CNN hybrid category, combining a transformer-based encoder with a convolutional decoder, enabling the model to capture global, non-local relationships more effectively than purely convolutional networks [32, 34]. This design allows the model to handle complex spatial patterns, which is especially important for tasks that require modeling long-range and cross variable dependencies.
Skip connections are a well-established component of encoder-decoder architectures, commonly used to preserve spatial detail and improve gradient flow. In ViT-CNN hybrid models, studies have shown that skip connections can enhance performance by bridging lower-level features from the encoder to the decoder [32]. However, in transformer-based architectures,especially those without hierarchical designs, this benefit may be less critical [33, 47]. Unlike CNN encoders, which typically reduce spatial resolution through downsampling, our ViT encoder preserves a constant resolution throughout. As a result, the output tokens retain the same spatial structure as the input, reducing the need for skip connections to recover lost detail.
@@
Specifically, our architecture is designed as follows: The encoding path follows a standard ViT configuration, where the input image is divided into non-overlapping $8\times 8$ patches, each linearly projected into a 256-dimensional vector by a dense embedding layer. To retain spatial information, learnable positional embeddings are added to the patch representations before passing them through the transformer stack. The encoder consists of six transformer layers, each following a consistent structure. Each layer applies normalization, followed by multi-head self-attention with eight heads and a dropout rate of 0.1. Next, A residual connection adds the input of the attention block to its output. After attention, a second normalization is applied, followed by a multi-layer perceptron (MLP) consisting of two fully connected layers with 512 and 256 units, respectively, and GELU [48] activation. Another residual connection is applied at the output of the MLP block. Through this process, the transformer encoder learns both local and global relationships between spatially distributed patches. To mitigate the memory constraints that limit batch size, we use Group Normalization [49] with a group size of 8, which has been shown to be more stable than Batch Normalization in such settings.
Following the transformer encoder, the sequence of embeddings is reshaped into a spatial grid, initializing the decoding path. The decoder is composed of three consecutive upsampling blocks, each of which includes bilinear interpolation for upsampling, doubling the spatial resolution at each step. We utilize bilinear interpolation avoiding artifacts often associated with learned transposed convolutions [50]. Each bilinear upsampling is followed by a residual block, composed of group normalization, two convolutional layers, and a residual connection that preserves input information. The first convolution in the block uses Swish activation, while the second uses no activation. The feature width begins at 256 in the first upsampling block and is halved in each subsequent block, reducing to 32 in the final block. A final convolutional layer with no activation and a kernel initializer set to zero projects the output to the target resolution, completing the decoding process. The complete architecture for a single encoder-decoder pair contains approximately 11 million learnable parameters.

### v2

@@ Abstract
Global Climate Models (GCMs) are critical for simulating large-scale climate dynamics, but their coarse spatial resolution limits their applicability in regional studies. Regional Climate Models (RCMs) address this limitation through dynamical downscaling, albeit at considerable computational cost and with limited flexibility. Deep learning has emerged as an efficient data-driven alternative; however, most existing approaches focus on single-variable models that downscale one variable at a time. This paradigm can lead to redundant computation, limited contextual awareness, and weak cross-variable interactions. To address these limitations, we propose a multi-variable Vision Transformer (ViT) architecture with a shared encoder and variable-specific decoders (1EMD). The proposed model jointly predicts six key climate variables: surface temperature, wind speed, 500 hPa geopotential height, total precipitation, surface downwelling shortwave radiation, and surface downwelling longwave radiation, directly from GCM-resolution inputs, emulating RCM-scale downscaling over Europe. Compared to single-variable ViT models, the 1EMD architecture improves performance across all six variables, achieving an average MSE reduction of approximately 5.5% under a fair and controlled comparison. It also consistently outperforms alternative multi-variable baselines, including a single-decoder ViT and a multi-variable U-Net. Moreover, multi-variable models substantially reduce computational cost, yielding a 29–32% lower inference time per variable compared to single-variable approaches. Overall, our results demonstrate that multi-variable modeling provides systematic advantages for high-resolution climate downscaling in terms of both accuracy and efficiency. Among the evaluated architectures, the proposed 1EMD ViT achieves the most favorable trade-off between predictive performance and computational cost.
@@ 1 Introduction
Global Climate Models (GCMs) are essential tools for simulating the Earth’s climate system and projecting future climate change. However, their coarse spatial resolution, typically around 1° (approximately 100 km, depending on the latitude), limits their ability to accurately capture regional-scale dynamics. To overcome this limitation, Regional Climate Models (RCMs) are employed to dynamically downscale GCM outputs, providing finer-scale climate information at resolutions typically ranging from 0.1° to 0.5° (respectively 10 to 50 km). Although RCMs enhance spatial detail and realism, they are computationally demanding, limiting the number of possible model runs in the context of regional coordinated experiments [Jacob2014].
@@
An emerging alternative to traditional numerical downscaling involves the use of deep learning techniques to emulate the RCM downscaling process. Instead of performing computationally intensive simulations, deep learning models learn to reproduce the mapping from coarse-resolution GCM outputs to high-resolution RCM fields from available data. This approach presents a rapid, data-driven method for producing detailed climate projections applicable across multiple regions and scenarios, increasing the number of available simulations [Rampal2024].
@@
In this study, we frame the GCM-to-RCM emulation task as a supervised deep learning problem, as depicted in Figure 1. Using paired datasets from GCM and RCM outputs, we employ Vision Transformer (ViT) models to learn the relationship between coarse-resolution inputs and their high-resolution counterparts over the European domain. The primary objective of this approach is to produce spatially detailed outputs that accurately capture the fine-scale structures inherent to RCM simulations, enabling efficient and cost-effective generation of high-resolution climate data.
Figure 1: Super-resolution problem setup. A low-resolution field (LR, left) is fed into a neural model (center), which outputs a super-resolved field (HR, right).
@@
The main contribution of this work is the exploration of joint, multi-variable processing. Typically, downscaling meteorological variables from GCM to RCM involves handling numerous variables spanning atmospheric and land-surface categories. Most existing deep learning approaches for climate downscaling focus exclusively on single-variable modeling, training and applying uni-variate models separately on each variable [bano2020configuration, vandal2017deepsd, oyama2023deep, sha2020deep, sha2020deep2]. This approach poses several issues: single-variable models cannot exploit valuable cross-variable interactions, managing and executing multiple independent models is inefficient, and similar meteorological variables result in redundant learning efforts when modeled separately. In response, this study explores multi-variable processing strategies.
@@
Figure 2: Visualization of a single daily sample for the six studied variables: Surface Temperature, Wind Speed, Geopotential Height, Total Precipitation, Surface Downwelling Shortwave Radiation and Surface Downwelling Longwave Radiation. Each variable is shown for both the high-resolution Regional Climate Model ICTP-RegCM4-6 (top row) and the low-resolution Global Climate Model MPI-ESM-LR (bottom row) re-gridded onto the same European domain.
@@
Vision Transformers (ViTs) have emerged as powerful architectures in image-to-image translation and super-resolution tasks due to their attention-based processing and ability to capture long-range dependencies [zhong2024investigating, yang2024successful, Rampal2024, liu2025spatial]. Leveraging this architecture, we propose processing several meteorological variables simultaneously, enabling the effective exploitation of inter-variable relationships and potentially enhancing overall performance and physical coherence. This approach significantly reduces the number of required models, thereby decreasing computational demands per variable and improving efficiency. However, multi-variable modeling also introduces challenges such as potential training instabilities and cross-variable interference, which may manifest as artifacts in outputs.
@@
To evaluate multi-variable modeling strategies for GCM-to-RCM downscaling, we leverage modern ViT architectures to model this image-to-image translation problem. Our study focuses on six atmospherical variables: wind speed (sfcWind), temperature (tas), geopotential height at 500 hPa (zg500), total precipitation (tp), surface downwelling shortwave radiation (rsds) and surface downwelling longwave radiation (rlds). An image depicting a sample of the six variables for both RCM and GCM is reported in Figure 2. We first establish baseline performance using single-variable ViT models trained independently for each variable. Subsequently, we introduce two multi-variable approaches: one treating variables as separate channels (1E1D), and a second employing a single shared encoder with multiple dedicated decoders (1EMD), one per variable. In addition, we incorporate a multi-variable U-Net model to provide an external and widely used baseline for comparison. We ensure a robust and reproducible comparison by standardizing model structure, parameter count, training epochs, learning rates, and optimizers across all experiments.
@@
Our findings reveal that multi-variable modeling offers clear advantages in terms of computational efficiency. However, the simpler multi-channel approach (1E1D) suffers from cross-variable interference, failing to match the performance of single-variable baselines. In contrast, our structured multi-decoder design (1EMD) avoids these issues and outperforms single-variable models across all tested variables. These results highlight the effectiveness of our multi-variable architecture for high-resolution climate downscaling.
@@ 1.1 Data Selection
Our study is based on the EURO-CORDEX data. EURO-CORDEX, developed as part of the Coordinated Regional Downscaling Experiment (CORDEX) under the World Climate Research Programme (WCRP), provides a consistent set of high-resolution (0.11°) RCM simulations, driven by boundary conditions from multiple GCMs and emission scenarios. These simulations are widely used for regional climate impact assessments across Europe [Jacob2014]. We subjectively selected one GCM-RCM couple from the available simulations: the MPI-ESM-LR GCM with the ICTP-RegCM4-6 RCM simulation under emissions scenario RCP8.5.
The selection of surface air temperature (tas), surface wind speed (sfcWind), geopotential height at 500 hPa (zg500), total precipitation (tp), surface downwelling shortwave radiation (rsds), and surface downwelling longwave radiation (rlds) was motivated by both practical relevance and methodological interest. These six variables exhibit markedly different spatial and statistical characteristics when downscaled. Temperature and geopotential height produce relatively smooth, large-scale fields, whereas wind speed contains localized high-frequency patterns. Precipitation, in contrast, is sparse, intermittent, and notoriously challenging to downscale. Radiative fluxes (rsds and rlds) reflect a mixture of smooth background structure and localized effects driven by cloud cover. This diversity offers a useful testbed for evaluating how well data-driven models handle different signal types, and whether modeling them jointly can facilitate positive transfer across variables.
The choice was also shaped by the broader goal of exploring multi-variable downscaling. By selecting a diverse set of variables, we can begin to assess whether downscaling performance benefits from shared representations across related but physically distinct fields. The downscaling task addresses a spatial resolution increase from 1.0° to 0.1°, representing a tenfold enhancement. This is among the largest resolution gaps available in publicly accessible GCM-RCM datasets, offering a particularly demanding test case for evaluating the capabilities of data-driven downscaling methods.
@@ 1.2 Related Works
In recent years, the focus of climate downscaling has increasingly shifted toward deep learning models [sun2024deep, bano2021suitability, bano2020configuration]. A wide range of architectures has been explored to downscale variables such as temperature, wind components, surface pressure, and precipitation, among others. These include convolutional neural networks (CNNs) [bano2021suitability, wang2021deep, dujardin2022wind, jin2022deep, damiani2024exploring, vandal2017deepsd], U-Nets [sha2020deep, hohlein2020comparative, doury2023regional], Vision Transformers (ViTs) [zhong2024investigating, yang2024successful, liu2025spatial], generative adversarial networks (GANs) [harris2022generative, kumar2023modern, stengel2020adversarial, manco2025new], and diffusion models [aich2024conditional, mardani2023generative, merizzi2024wind, merizzi2025controlling, watt2024generative].
Different architectures show varying strengths depending on the target variable. Generative models like GANs and diffusion models are particularly effective at preserving high-frequency spatial detail and have shown strong performance in downscaling wind speed [tomasi2025can]. In contrast, smoother fields such as temperature are often better handled by U-Nets [liu2024enhanced]. Despite these findings, there is no clear consensus on a universally optimal architecture. Recently, ViT-based models have shown competitive and often superior performance across multiple climate variables, outperforming CNN-based baselines in several studies [loganathan2025regional, chen2022rainnet, zhong2024investigating]. As such, ViTs form the foundation of our proposed approach.
ViT-based architectures for image-to-image tasks often draw inspiration from the encoder-decoder structure of U-Nets. One common design integrates a transformer-based encoder with a convolutional decoder [chen2021transunet, ranftl2021vision, liang2021swinir], while other variants opt for fully transformer-based pipelines [wang2022uformer, zheng2022ittr]. A key architectural consideration is the treatment of skip connections, which preserve spatial resolution and improve gradient flow. These may be preserved directly [chen2021transunet, zhang2025alternate], omitted entirely under the assumption of rich encoder representations [ranftl2021vision], or implemented through attention-based gating mechanisms [wang2022uformer].
@@
Multi-variable downscaling naturally aligns with the framework of multi-task learning, where a single model learns to solve multiple related tasks simultaneously. In climate applications, multi-task learning offers several potential advantages: improved generalization through shared representations, reduced training and inference cost, and the possibility of positive transfer across variables [crawshaw2020multi].
Several recent studies support the promise of this approach [schillinger2025enscale]. For instance, Ling et al. [ling2022multi] present a multi-task model for seasonal climate prediction that jointly forecasts several indices related to the Indian Ocean Dipole. The model is trained using a loss function defined as a weighted sum of the individual task with a higher loss weight to the primary target (DMI). The model achieves improved performance compared to single-variable baselines. Similarly, Fan et al. [Fan2020] propose M2GSNet, a graph-based multi-task model for wind power prediction across multiple farms. By utilizing shared encoders across tasks and training with an unweighted sum of L1 losses across farms, the model achieves superior forecasting accuracy compared to single-variable baselines.
Other examples further support the benefits of multi-task formulations. Bannai et al. [bannai2023multi] use a shared CNN backbone with two heads for simultaneous rain/no-rain classification and rain rate estimation, yielding improved performance through joint optimization of a combined loss function. Liu et al. [liu2023dsaf] propose a two-stage downscaling system incorporating bias correction, terrain inputs, and a multi-task loss with physical constraints, demonstrating the potential of multi-variable models for enhancing both accuracy and realism in climate data products. Quesada-Chacón et al. [quesada2023downscaling] propose a deep learning-based downscaling framework that employs U-Net architectures to generate daily 1km resolution gridded ensembles of seven meteorological variables, including precipitation, temperature (min/mean/max), radiation, humidity, and wind.
@@ 2 Methodology
Figure 3: Comparison between a single-variable encoder-decoder model and multi-variable models using either a single shared decoder (1E1D) or multiple decoders branching from a shared encoder (1EMD).
@@ 2.1 Architectural Backbone: Vision Transformer (ViT)
In recent years ViTs have been successfully applied to a variety of super-resolution tasks [lu2022transformer, 10836746, cao2021video, liu2023image]. Our proposed ViT architecture belongs to the ViT-CNN hybrid category, combining a transformer-based encoder with a convolutional decoder, enabling the model to capture global, non-local relationships more effectively than purely convolutional networks [chen2021transunet, liang2021swinir]. This design allows the model to handle complex spatial patterns, which is especially important for tasks that require modeling long-range and cross-variable dependencies.
Skip connections are a well-established component of encoder-decoder architectures, commonly used to preserve spatial detail and improve gradient flow. In ViT-CNN hybrid models, studies have shown that skip connections can enhance performance by bridging lower-level features from the encoder to the decoder [chen2021transunet]. However, in transformer-based architectures,especially those without hierarchical designs, this benefit may be less critical [ranftl2021vision, zheng2021rethinking]. Unlike CNN encoders, which typically reduce spatial resolution through downsampling, our ViT encoder preserves a constant resolution throughout. As a result, the output tokens retain the same spatial structure as the input, reducing the need for skip connections to recover lost detail.
@@
Specifically, our architecture is designed as follows: The encoding path follows a standard ViT configuration, where the input image is divided into non-overlapping $8\times 8$ patches, each linearly projected into a 256-dimensional vector by a dense embedding layer. To retain spatial information, learnable positional embeddings are added to the patch representations before passing them through the transformer stack. The encoder consists of six transformer layers, each following a consistent structure. Each layer applies normalization, followed by multi-head self-attention with six heads and a dropout rate of 0.1. Next, A residual connection adds the input of the attention block to its output. After attention, a second normalization is applied, followed by a multi-layer perceptron (MLP) consisting of two fully connected layers with 512 and 256 units, respectively, and GELU [hendrycks2016gaussian] activation. Another residual connection is applied at the output of the MLP block. Through this process, the transformer encoder learns both local and global relationships between spatially distributed patches. To mitigate the memory constraints that limit batch size, we use Group Normalization [Wu2018] with a group size of 8, which has been shown to be more stable than Batch Normalization in such settings.
Following the transformer encoder, the sequence of embeddings is reshaped into a spatial grid, initializing the decoding path. The decoder is composed of three consecutive upsampling blocks, each of which includes bilinear interpolation for upsampling, doubling the spatial resolution at each step. We utilize bilinear interpolation avoiding artifacts often associated with learned transposed convolutions [odena2016deconvolution]. Each bilinear upsampling is followed by a residual block, composed of group normalization, two convolutional layers, and a residual connection that preserves input information. The first convolution in the block uses Swish activation, while the second uses no activation. The feature width begins at 256 in the first upsampling block and is halved in each subsequent block, reducing to 32 in the final block. A final convolutional layer with no activation and a kernel initializer set to zero projects the output to the target resolution, completing the decoding process. The complete architecture for a single encoder-decoder pair contains approximately 11 million learnable parameters.